# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ 7.3 Baselines
As baselines, we consider the performance of two popular state-of-the-art algorithms. One is the Best-of-64 (snell2024scaling) baseline, which is one of the most popular and well-performing test-time alignment methods that has been reported to outperform reinforcement learning fine-tuning methods under large enough $N$ (gao2023scaling; mudgal2023controlled; eisenstein2023helping; beirami2024theoretical). Specifically, the test persona set-based evaluation result for the Best-of-64 ad (identified above in Section 7.2 using training persona set) was used as the Best-of-64 baseline. We also evaluated the ads in Worst5-of-64 in the same way and took the mean of their scores to obtain the Worst5-of-64 baseline.
Another baseline is GEPA (Genetic-Pareto) (agrawal2025gepa). As discussed earlier, this reflective and evolutionary prompt evolution algorithm has recently demonstrated superior performance compared to GRPO (liu2024deepseek), a state-of-the-art reinforcement fine-tuning technique. GEPA keeps a set of candidate prompts, and samples a prompt to mutate based on its evaluation result for a problem (in our case, a persona) set they call the Pareto set. A Pareto set stores ‘important’ problems. Since we randomly sample personas from the training persona set, we cannot store specific personas and retrieve them later. Therefore, we instead use the version of GEPA that 1) randomly samples a persona and 2) mutates all candidate prompts equally and throws away the old prompts to avoid unbounded growth of the candidate set.
For the experiment, we run five-trajectory ($J=5$) parallel T-BoN BO and GEPA starting from Worst5-of-64, i.e., the worst five initial prompts we identified in Section 7.2. This starting point allows us to attribute observed gains to the algorithms rather than to favorable starting prompts. We test if T-BoN BO outperforms the Best-of-N baseline and GEPA baseline throughout the optimization process. For both T-BoN BO and GEPA, we ran 10 optimization steps, evaluating 200 randomly sampled personas per step (2,000 total persona evaluations). We used a multi-modal LLM (Gemini-2.5-Flash) to serve as the critic model for both T-BoN BO and GEPA.
@@ 7.4 Results and Analysis
Trend analysis
Figure 8: Progress comparison of T-BoN BO and GEPA with Best-of-64 baseline. T-BoN BO implements parallel T-BoN BO with $J=5$. The performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set. Since it is the mean of means, we don’t plot the standard errors.
Figure 8 plots how much T-BoN BO, GEPA, and Best-of-N improved their performance from the initial Worst5-of-N score across eight scenarios. (For the detailed data used to generate this plot, refer to Table 2 and 3 in Appendix C.) At each step (the $x$-axis), the performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set. For example, the step-2 score is the average of the eight step-2 mean evaluation scores. As the figure illustrates, both T-BoN BO and GEPA improve consistently across steps, with T-BoN BO demonstrating faster early gains and achieving a significantly better final performance than GEPA. Notably, by the third step on average, which is equivalent to 600 persona evaluations, T-BoN BO and GEPA both outperform the Best-of-64 baseline.
Statistical analysis
Model
$\text{score}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TBoN}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon$
@@
$\beta_{2}$: T-BoN BO gain - GEPA gain $(t=10)$ from Worst5-of-64 0.0743*** 0.0248 2.991
$\beta_{1}+\beta_{2}$: T-BoN BO gain $(t=10)$ from Worst5-of-64 0.3911*** 0.0355 11.011
@@
Table 1: Two-way fixed effect (TWFE) analysis result with persona and scenario fixed effects with two-way clustered SEs (cameron2011robust). Standard error of $\beta_{1}+\beta_{2}$ was computed using Delta method.
In addition to the aggregate trend analysis provided in Figure 8, we now provide statistical evidence that T-BoN BO indeed outperforms the GEPA. For this, we conducted a Two-way Fixed Effect (TWFE) analysis using T-BoN BO and GEPA’s final ($t=10$) scores to compare the effectiveness of each method rigorously. Specifically, we consider the following TWFE model:
@@
$\text{score}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TBoN}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon.$ (27)
where $\mathbf{1}\{t=10\}$ is the indicator variable of whether it is $t=0$ (GEPA and T-BoN BO’s initial starting point, i.e., Worst5-of-64 score) or $t=10$ (GEPA and T-BoN BO’s final score), $\beta_{1}$ is the GEPA’s gain from Worst5-of-64 after $t=10$, $\beta_{2}$ is the T-BoN BO’s advantage over GEPA after $t=10$, and $\beta_{1}+\beta_{2}$ is the T-BoN BO’s gain from Worst5-of-64 after $t=10$. The unobserved residual $\varepsilon$ is assumed to follow the normal distribution. To remove potential correlation among unobserved residuals within the same persona or scenario, we conducted a cluster-robust Two-way Fixed Effect (TWFE) analysis with two-way clustering (cameron2011robust; cameron2015practitioner). The result of TWFE analysis is shown in Table 1. As we can see, both of T-BoN BO’s gain and GEPA ’s gain from the starting point (Worst5-of-64) after $T=10$ are statistically significant. In addition, we conclude that T-BoN BO has a significant 23.5% more gain in score compared to GEPA’s gain.
Qualitative analysis
The significant gains of GEPA and T-BoN BO can be interpreted as demonstrating their ability to align with the persona distribution. However, such significance can also be easily gained by starting from a poor starting point, i.e., a poor Worst5-of-N prompt and corresponding ad image. One way to check whether this happens is to compare the quality of the prompts; we provide them in Appendix A to demonstrate that this is not the case. Another qualitative approach is to compare the generated ad images and verify whether the differences among ads are in quality or in the message and representation. Figure 9 compares the ad images of the baseline methods. As we can see, there is no clear visual quality difference among those ad images; the differences are mainly in the ad’s message and how it is represented. This supports the claim that T-BoN BO and GEPA improvements from Worst5-of-N demonstrated in Figure 8 are attributed to alignment, not trivial quality gains from correcting poor or under-specified prompts.
@@
T-BoN BO (after T=10) Refer to caption Score: 3.168 Refer to caption Score: 3.132 Refer to caption Score: 3.104
Figure 9: T-BoN BO and baseline’s generated Ad images for three fictional brands: GreenBite burger patty (Scenario 1), MindGarden meditation app (Scenario 6), and AuraSonics noise-canceling earbuds (Scenario 2). For the Ad images of the other five brand scenarios, see Figure 17 in Appendix C.
Figure 10: An ablation study experimenting T-BoN BO with $J=5$ and Best-of-64 for Gemini 2.5 Flash instead of persona distribution with temperature 0. (GEPA was excluded, as there is no problem-dependent heterogeneity.) The performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set. Again, since it is the mean of means, we don’t plot the standard deviations.
@@ 7.5 Ablation study
The main experiment in Section 7.1-7.4 evaluates the candidate ad at each iteration as the mean of scores across 411 personas (20% of the persona data in Twin-2k-500 (toubia2025twin)) that were injected into the LLM’s context. Each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens. Would T-BoN BO still outperform strong state-of-the-art baselines when the LLM receives less persona context information?
@@
To test this, we ran a no-persona ablation experiment in which the evaluator is a single LLM (Gemini 2.5 Flash) with no auxiliary persona context. Concretely, at each iteration, we present only the ad image candidate to LLM judge (Gemini 2.5 Flash) together with the same 1–5 effectiveness rubric, set temperature to 0, and compute the same scalar score as the log-probability–weighted expectation over $\{1,\dots,5\}$ as before. [^11] With temperature =0=0, the judge’s output is deterministic for a fixed image and rubric; the logprob vector is still available, so the expected score is stable across repeated calls. This collapses the evaluation to a deterministic mapping and removes heterogeneity from personas and sampling, allowing us to test whether T-BoN BO’s performance depends on specific persona information-related setup. To ensure there is no leakage of preference, we used GPT-5 instead of Gemini 2.5 Flash for generating meta-reflection, generating textual gradients, and making Best-of-N choices.
We again compare T-BoN BO (J=5J{=}5 trajectories) against the Best-of-64 initial baseline; we exclude GEPA, which requires multiple personas (multiple problems, in GEPA’s language.)
Figure 10 plots T-BoN BO across 10 optimization steps starting from the Worst5-of-64 prompts under the ablation setup. As in Figure 8, the main experiment’s results, the value shown in Figure 10 is the mean of that step’s scores over eight scenarios. As can see in Figure 10, T-BoN BO again outperforms Best-of-64 baseline from optimization step 3. Except for the fact that in general the evaluation scores without persona information are more optimistic than scores with persona information for both T-BoN BO and Best-of-64, the ablation experiment’s trend in Figure 10 shows the same trend we saw in Figure 8.

### v3

@@ 8.1 Baselines
We consider two state-of-the-art algorithms in the AI literature as baselines – (1) Best-of-N (snell2024scaling) and (2) GEPA (agrawal2025gepa).
Best-of-N is one of the most popular and well-performing test-time alignment methods that has been reported to outperform reinforcement learning fine-tuning methods under large enough $N$ (gao2023scaling; mudgal2023controlled; eisenstein2023helping; beirami2024theoretical). For any given scenario, this algorithm samples $N$ candidate outputs from a fixed generator and returns the single candidate with the highest predicted score under a judge/reward model, trading extra inference-time compute for higher expected quality (snell2024scaling; beirami2024theoretical). In our setting, recall that we generate a set of 64 initial ads (see $\S$7.2 on the details of the initial ad set generation). Then, we use the training dataset of personas and the Eval() operator in conjunction witha best-arm identification algorithm to identify the Best-of-64 ad (as described in $\S$7.3). Finally, we evaluate Best-of-64’s performance on the testing persona dataset; this score serves as our Best-of-64 baseline for that scenario.
GEPA is a reflective and evolutionary prompt optimization algorithm that has demonstrated superior performance compared to state-of-the-art reinforcement fine-tuning techniques such as GRPO (liu2024deepseek). In our implementation of GEPA, we use the same set of initial prompts, critic models, ad generation modules, and evaluation operators we use for TextBO (see $\S$7.2 for details). Further, for each of the eight scenarios, we start GEPA with the same Worst5-of-64 prompts as TextBO, i.e., the worst five initial prompts we identified in $\S$7.2. We also run GEPA for the same number of iterations ($T=10$). Here, keeping the components of the algorithms and the starting points the same across both algorithms ensures a fair comparison. Notably, when we adapt GEPA to the digital ad optimization problem setup (where we cannot store and retrieve problem instances (e.g., personas) freely anytime we want as in agentic AI benchmarks of agrawal2025gepa), it effectively reduces to TextBO with $G=1$ and N=1N=1121212In $\S$9, we adapt the principles of Best-of-N gradient multiple gradient steps of TextB for GEPA and test it for agentic AI benchmark experiments (agrawal2025gepa) for which GEPA was tested.; see Web Appendix D.1 for the implementation details of the adapted GEPA.
8.2 Experiment Design
Our experiment is designed to test whether we can start from ex-post least-aligned ads (Worst5-of-64) and apply TextBO to outperform the ex-post best-aligned ad (Best-of-64). Here, “alignment” is measured by the evaluation score induced by our persona-based evaluation environment (see $\S$7.2). This design choice matters because the Best-of-64 baseline is already a strong, ex-post selection benchmark: it is the best-performing ad within a reasonably rich, human-like initial pool of 64 creatives generated from the same creative brief (snell2024scaling; beirami2024theoretical). By initializing optimization from the opposite end of the same pool (Worst5-of-64), we attribute improvements to the algorithm’s performance rather than to a favorable starting prompt. Relatedly, we are also interested in comparing the performance of TextBO with GEPA. To do so, we also need to ensure that the comparison between GEPA and TextBO is fair and both use similar starting points and algorithm components. Therefore, for each scenario, we proceed as follows:
  • $\bullet$
    Construct a common initial candidate pool. Using the scenario’s creative brief, we generate 64 detailed prompts and their corresponding ad images. All algorithms start with this initial pool of prompts and ads.
  • $\bullet$
    Define ex-post baselines and initializations within the pool. Using the training persona set and the fixed evaluation operator $\mathrm{Eval}(\cdot)$, we identify (i) the single best ad in the pool (Best-of-64) and (ii) the five worst ads (Worst5-of-64) via best-arm identification (see $\S$7.3). We use Worst5-of-64 as the starting set $\{\pi_{0}^{j}\}_{j=1}^{5}$ for both TextBO and GEPA, while Best-of-64 serves as the ex-post “best-aligned” benchmark.
  • $\bullet$
    Hold the algorithm components and evaluation environment fixed across methods. All methods use the same (i) scenarios and initial pool, (ii) ad generator $\Phi$, (iii) critic model $M_{\text{critic}}$, and (iv) persona-based evaluation operator. This ensures that differences in outcomes are attributable to the optimization procedure, rather than the evaluation environment.
  • $\bullet$
    Evaluate under a common budget and report out-of-sample performance. During optimization steps in GEPA and TextBO, each costly evaluation uses the same protocol (randomly sampled 200 personas from the training set; see $\S$7.2). For reporting, we evaluate the resulting ads on the held-out 411 test persona set and compare progress trajectories and final performance across methods.
@@ 8.3 Results and Analysis
We now present the main results from the experiments. Note that while the algorithm is designed to minimize regret, we cannot directly plot regret here since we don’t have true $\pi^{\star}$ (or the the prompt that maximizes the score), a priori, for any of the scenarios. Therefore, throughout, we report experiment results in terms of empirical performance scores.
Figure 5: Progress comparison of TextBO and GEPA with Best-of-64 baseline. TextBO implements parallel TextBO with $J=5$. The performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the testing persona set.
Main results and trend analysis:
Figure 5 shows the results of our numerical experiments over 10 iterations: it plots how much TextBO and GEPA improved their performance from the initial Worst5-of-N score across eight scenarios. (For the detailed data used to generate this plot, see Tables A1 and A2 in Web Appendix D.2.) At each step (the $x$-axis), the performance score (the $y$-axis) is the average of eight scenarios’ mean evaluation score for the test persona dataset. [^13] Best-of-N remains constant because, as discussed earlier, it is simply the best among the initial pool of ads and this baseline does not involve generation of new ads. For example, the step 2 score is the average of the eight step 2 mean evaluation scores. Both TextBO and GEPA improve consistently across steps, with TextBO demonstrating faster early gains and achieving a significantly better final performance than GEPA. Notably, by the third step on average, which is equivalent to 600 persona evaluations, both TextBO and GEPA outperform the Best-of-64 baseline. Further, by step 8, both algorithms have reached stable performance.
@@
$\beta_{2}$: TextBO gain - GEPA gain $(t=10)$ from Worst5-of-64 0.0743*** 0.0248 2.991
$\beta_{1}+\beta_{2}$: TextBO gain $(t=10)$ from Worst5-of-64 0.3911*** 0.0355 11.011
@@
Table 1: Two-way fixed effect (TWFE) analysis result with persona and scenario fixed effects with two-way clustered SEs. Standard error of $\beta_{1}+\beta_{2}$ is computed using the Delta method.
Statistical analysis:
Figure 5 shows the mean of means (i.e., scores are averaged over personas in test dataset and then over the eight scenarios). Since both the set of personas and scenarios are common across all algorithms, we cannot directly plot standard errors. Therefore, we now present a statistical analysis using the raw scores at the persona-scenario level and examine whether the performance of TextBO is statistically significant different from that of GEPA. We consider the Two-way Fixed Effect (TWFE) model:
@@
$\text{score(persona, scenario, t)}=\alpha+\beta_{1}\mathbf{1}\{t=10\}+\beta_{2}\mathbf{1}\{\mathrm{TextBO}\}\,\mathbf{1}\{t=10\}+\text{persona FE}+\text{scenario FE}+\varepsilon,$
where $\mathbf{1}\{t=10\}$ is the indicator variable of whether it is $t=0$ (GEPA and TextBO’s initial starting point, i.e., Worst5-of-64 score) or $t=10$ (GEPA and TextBO’s final score), $\beta_{1}$ is the GEPA’s gain from Worst5-of-64 after $t=10$, $\beta_{2}$ is the TextBO’s advantage over GEPA after $t=10$, and $\beta_{1}+\beta_{2}$ is the TextBO’s gain from Worst5-of-64 after $t=10$. Further, we control for persona and scenario fixed effects to account for persistent differences in personas and scenarios (across both algorithms). The standard errors are two-way clustered (cameron2011robust; cameron2015practitioner).
The results from this TWFE model are shown in Table 1. We see that both TextBO’s and GEPA show statistically significant gains from the starting point (Worst5-of-64) after ten iterations. In addition, at $T=10$, TextBO shows a 23.5% more gain in score compared to GEPA. Together, these findings provide two key takeaways: (1) both GEPA and TextBO are able to self-improve and generate ads better aligned with the target population, and (2) the performance of TextBO is significantly better than that of GEPA for the same number of evaluations, and this difference is both statistically significant and meaningful in magnitude.
Qualitative analysis:
So far, we have shown that TextBO is able to self-improve and generate ads that align well with the persona distribution within a few iterations. However, one could argue that this can be simply due to a weak set of initial prompts, i.e., a poor set of Worst5-of-N prompts and corresponding ad images. We examine this alternative explanation by examining the prompts and ad creatives/images.
First, one natural way to rule out this explanation is to compare the quality of the prompts at the starting point and end point of the algorithm, and show that the prompts at the starting point are grammatically correct, meaningful, and sensible and that the differences in the prompts at $t=0$ and $t=10$ stem from differences in the message and focus of the prompt rather than grammar/structure. We confirm that this in indeed the case by comparing the prompts at the starting and end points of the algorithm for all the scenarios and trajectories. For example, see the prompt for the Worst-of-64 prompt for Scenario 1 in Figure A2 in Web Appendix C.2 and the prompt for TextBO at the last step in Figure A9 in Web Appendix D.3.
A second way to rule this out this alternative explanation is to compare the generated ad images and verify that the differences among ads across algorithms and (over the iterations, within a algorithm) stem from the message and representation, and not from the image quality. To that end, in Figure 6, we show the ad images at the start and end points for both GEPA and TextBO, as well the Best-of-N baseline (for three scenarios). There is no clear visual quality differences among these ad images; the differences are mainly in the ads’ message and how it is represented. This further suggests that the improvements from TextBO (and GEPA) over the Worst5-of-N starting point can be attributed to alignment, and are not due to trivial quality gains from correcting poor or under-specified initial prompts.
@@
TextBO (after T=10) Refer to caption Score: 3.168 Refer to caption Score: 3.132 Refer to caption Score: 3.104
Figure 6: TextBO and baselines’ generated ad images for three fictional brands: GreenBite burger patty (Scenario 1), MindGarden meditation app (Scenario 6), and AuraSonics noise-canceling earbuds (Scenario 2). For the ad images of the other five scenarios, see Figure A8 in Web Appendix D.2.
Figure 7: Ablation study with TextBO with $J=5$ and Best-of-64 for Gemini 2.5 Flash as the evaluator, with temperature 0. The performance score is the average of eight scenarios’ mean evaluation score at that step.
@@ 8.4 Ablation Study
So far, in all our experiments, we used the persona dataset for evaluation and the goal was to optimize the ads serve to the target population captured by the persona dataset (toubia2025twin). However, a natural question here is whether TextBO can still perform well if we simply use a different target population/evaluator without access to extensive persona information during training. (Recall that each persona consists of a large amount of information, with 500 survey questions accounting for around 50,00050{,}000 tokens.) Therefore, we run a no-persona ablation experiment in which the evaluator is a single LLM (Gemini 2.5 Flash) with no auxiliary persona context. Concretely, at each iteration, we present only the ad image candidate to a LLM judge (Gemini 2.5 Flash) together with the same 1–5 effectiveness rubric, set the temperature to 0, and compute the same scalar score as the log-probability–weighted expectation over $\{1,\dots,5\}$ as before. [^14] With temperature =0=0, the judge’s output is deterministic for a fixed image and rubric; the logprob vector is still available, so the expected score is stable across repeated calls. This collapses the evaluation to a deterministic mapping between ads and scores; it removes heterogeneity from personas and stochasticity from sampling, allowing us to test whether TextBO’s performance depends on specific persona information-related setup. To ensure that there is no leakage of preferences, we used GPT-5 instead of Gemini 2.5 Flash for critic model $M_{critic}$ (i.e., the operators used for generating meta-reflection, generating textual gradients, and making the Best-of-N choices.).
We again compare TextBO (J=5J{=}5 trajectories) against the Best-of-64 initial baseline. The results from this exercise are shown in Figure 7 plots TextBO across 10 optimization steps starting from the Worst5-of-64 prompts under the ablation setup. [^15] Here, Worst5-of-64 and Best-of-64 are different from those of the main experiments, as the preference distribution is different. As before, we see that TextBO outperforms Best-of-64 baseline from optimization step 3 onwards, and the overall trend in Figure 7 is similar to that in Figure 5. The main difference is that the evaluation scores without persona information are more optimistic than scores with persona information for both TextBO and Best-of-64. This is understandable, since peng2025mega find that persona information does not significantly improve LLMs’ overall ability to predict preferences and mainly adds variance to the predictions. Since variance makes the learning problem harder, it is natural that the main experiment, which uses personas, shows slower improvement than the pure-LLM case here. Overall, these findings suggest that when learning the target population’s preferences is difficult, the method may require more iterations and larger training samples to optimize effectively.
9 Experiments for Agentic AI Benchmarks
In $\S$8, we saw that TextBO significantly outperforms GEPA for ad optimization experiments. Recall that the adapted version of GEPA in the ads settings effectively reduced to TextBO with $N=1$ and $G=1$. As such, one of the main takeaways from those experiments is that Best-of-N gradient sampling and multiple gradient steps can effectively improve the performance of self-improving AI under limited evaluation budget.
[…]
9.1 Experiment Design
To rigorously evaluate GEPA and TextBO-GEPA, we closely replicate three agentic AI benchmarks and their corresponding experimental setups from agrawal2025gepa.161616agrawal2025gepa also included IFBench ifbench_bench, for which they report that none of the algorithm show improvement. Therefore we exclude it in our analysis. In all the experiments, we used Qwen3-8B (qwen3_technical_report), the same model and setup used in agrawal2025gepa (decoding temperature of 0.6, top-p of 0.95, and top-k of 20 for training as well as inference). As reported in agrawal2025gepa, GEPA saturates after $7,500$ evaluations (rollouts), so we conduct experiments up to $7,500$ evaluations.
We now briefly describe the three benchmark tasks:
HotpotQAhotpotqa_bench provides HotpotQA, a set of diverse, explainable multi-hop question answering tasks comprising approximately 113k Wikipedia-based examples. This task requires reasoning over multiple supporting documents and providing sentence-level supporting facts to enable intense supervision. We follow the implementation details described in agrawal2025gepa, except for the retriever module, which was not specified in the paper. So we conservatively set it to BM25 (robertson2009probabilistic), one of the most widely used retrievers. We use 150 examples for training, 300 for validation, and 300 for testing.
HoVerhover_bench provides HoVer, a set of many-hop fact-extraction and claim-verification tasks, constructed from the same 113k Wikipedia-based examples used in HotpotQA. It tests complex reasoning that requires aggregating evidence from multiple sentences across disparate documents, often involving diverse reasoning graphs and long-range dependencies. Again, we follow the implementation details described in agrawal2025gepa, except for the retriever module, which was again not specified in the paper, and was set to BM25 (robertson2009probabilistic). Again, we use 150 examples for training, 300 for validation, and 300 for testing.
PUPApapillon_bench provides PUPA, a set of privacy-conscious delegation tasks which address real-world user queries by orchestrating an ensemble of trusted and untrusted models. The primary objective is to maintain a high response quality, comparable to that of proprietary models, while minimizing exposure of personally identifiable information (PII) to the untrusted backend. Again, we follow the implementation details described in agrawal2025gepa, except that we use case-insensitive substring matching instead of LLM-as-a-judge to check whether each PII exists in the output. We use 111 training examples, 111 for validation, and 221 for testing.
9.2 Results and Analysis
(a) HotpotQA, Qwen3 8B
(b) HoVer, Qwen3 8B
(c) PUPA, Qwen3 8B
Figure 8: Experiment results for three agentic AI benchmarks considered in agrawal2025gepa. As more evaluations are conducted, TextBO-GEPA (GEPA with Best-of-N gradient sampling and multiple gradient steps) consistently outperform vanilla GEPA.
Figure 8 and Table 2 present the results of agentic AI benchmark experiments. We plot the improvement from their initial scores for both TextBO-GPEA and GEPA. Relative to the baseline, TextBO-GEPA delivers larger gains than GEPA across all datasets. On HotpotQA, the improvement rises from +1.00 to +3.00 points, which is a $200\%$ increase in the baseline-relative gain. On Hover, the gain increases from +5.00 to +7.67 points $(+53.4\%)$. On PUPA, it grows from +10.28 to +15.95 points $(+55.2\%)$. Overall, the aggregate improvement increases from +5.43 to +8.87 points, corresponding to a $63.4\%$ increase in baseline-relative improvement. Considering that language model evaluation benchmarks often have standard deviations well below 1 point, and 95% CI sizes often around 1–2 points (madaan2024quantifying),these results illustrate that TextBO-GEPA consistently significantly outperforms GEPA across the agentic AI benchmarks.
Model HotpotQA Hover PUPA Aggregate Improvement
Qwen3-8B
Baseline 56.67 41.33 66.58 54.86
GEPA 57.67 46.33 76.86 60.29 +5.43
TextBO-GEPA 59.67 49.00 82.53 63.73 +8.87
Table 2: Results on HotpotQA, Hover, and PUPA for Qwen3-8B after  7,500 evaluations. Aggregate is the mean across the three datasets; Improvement is vs. Baseline aggregate.